### v1

# Multistep Consistency Models

Abstract
Diffusion models are relatively easy to train but require many steps to generate samples. Consistency models are far more difficult to train, but generate samples in a single step.
In this paper we propose Multistep Consistency Models: A unification between Consistency Models (Song et al., 2023) and TRACT (Berthelot et al., 2023) that can interpolate between a consistency model and a diffusion model: a trade-off between sampling speed and sampling quality. Specifically, a 1-step consistency model is a conventional consistency model whereas we show that a $\infty$-step consistency model is a diffusion model.
Multistep Consistency Models work really well in practice. By increasing the sample budget from a single step to 2-8 steps, we can train models more easily that generate higher quality samples, while retaining much of the sampling speed benefits. Notable results are 1.4 FID on Imagenet 64 in 8 step and 2.1 FID on Imagenet128 in 8 steps with consistency distillation. We also show that our method scales to a text-to-image diffusion model, generating samples that are very close to the quality of the original model.
Machine Learning, ICML
1 Introduction
Diffusion models have rapidly become one of the dominant generative models for image, video and audio generation (Ho et al., 2020; Kong et al., 2021; Saharia et al., 2022). The biggest downside to diffusion models is their relatively expensive sampling procedure: whereas training uses a single function evaluation per datapoint, it requires many (sometimes hundreds) of evaluations to generate a sample.
Recently, Consistency Models (Song et al., 2023) have reduced sampling time significantly, but at the expense of image quality. Consistency models come in two variants: Consistency Training (CT) and Consistency Distillation (CD) and both have considerably improved performance compared to earlier works. TRACT (Berthelot et al., 2023) focuses solely on distillation with an approach similar to consistency distillation, and shows that dividing the diffusion trajectory in stages can improve performance. Despite their successes, neither of these works attain performance close to a standard diffusion baseline.
Here, we propose a unification of Consistency Models and TRACT, that closes the performance gap between standard diffusion performance and low-step variants. We relax the single-step constraint from consistency models to allow ourselves as much as 4, 8 or 16 function evaluations for certain settings. Further, we generalize TRACT to consistency training and adapt step schedule annealing and synchronized dropout from consistency modelling. We also show that as steps increase, Multistep CT becomes a diffusion model. We introduce a unifying training algorithm to train what we call Multistep Consistency Models, which splits the diffusion process from data to noise into predefined segments. For each segment a separate consistency model is trained, while sharing the same parameters. For both CT and CD, this turns out to be easier to model and leads to significantly improved performance with fewer steps. Surprisingly, we can perfectly match baseline diffusion model performance with only eight steps, on both Imagenet64 and Imagenet128.
Another important contribution of this paper that makes the previous result possible, is a deterministic sampler for diffusion models that can obtain competitive performance on more complicated datasets such as ImageNet128 in terms of FID score. We name this sampler Adjusted DDIM (aDDIM), which essentially inflates the noise prediction to correct for the integration error that produces blurrier samples.
In terms of numbers, we achieve performance rivalling standard diffusion approaches with as little as 8 and sometimes 4 sampling steps. These impressive results are both for consistency training and distillation. A remarkable result is that with only 4 sampling steps, multistep consistency models obtain performances of 1.6 FID on ImageNet64 and 2.3 FID on Imagenet128.
Figure 1: This figure shows that Multistep Consistency Models interpolate between (single step) Consistency Models and standard diffusion. Top at $t=0$: the data distribution which is a mixture of two normal distributions. Bottom at $t=1$: standard normal distribution. Left to right: the sampling trajectories of (1, 2, 4, $\infty$)-step Consistency Models (the latter is in fact a standard diffusion with DDIM) are shown. The visualized trajectories are real from trained Multistep Consistency Models. The 4-step path has a smoother path and will likely be easier to learn than the 1-step path.
2 Background: Diffusion Models
Diffusion models are specified by a destruction process that adds noise to destroy data: ${\bm{z}}_{t}=\alpha_{t}{\bm{x}}+\sigma_{t}{\bm{{\epsilon}}}_{t}$ where ${\bm{{\epsilon}}}_{t}\sim\mathcal{N}(0,1)$. Typically for $t\to 1$, ${\bm{z}}_{t}$ is approximately distributed as a standard normal and for $t\to 0$ it is approximately ${\bm{x}}$. In terms of distributions one can write the diffusion process as:
$q({\bm{z}}_{t}|{\bm{x}})=\mathcal{N}({\bm{z}}_{t}|\alpha_{t}{\bm{x}},\sigma_{t}).$ (1)
Following (Sohl-Dickstein et al., 2015; Ho et al., 2020) we will let $\sigma_{t}^{2}=1-\alpha_{t}^{2}$ (variance preserving). As shown in (Kingma et al., 2021), the specific values of $\sigma_{t}$ and $\alpha_{t}$ do not really matter. Whether the process is variance preserving or exploding or something else, they can always be re-parameterized into the other form. Instead, it is their ratio that matters and thus it can be helpful to define the signal-to-noise ratio, i.e. $\mathrm{SNR}(t)=\alpha_{t}^{2}/\sigma_{t}^{2}$.
To sample from these models, one uses the denoising equation:
$q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{x}})=\mathcal{N}({\bm{z}}_{s}|\mu_{t\to s}({\bm{z}}_{t},{\bm{x}}),\sigma_{t\to s})$ (2)
where ${\bm{x}}$ is approximated via a learned function that predicts $\hat{{\bm{x}}}=f({\bm{z}}_{t},t)$. Note here that $\sigma_{t\to s}^{2}=\big{(}\frac{1}{\sigma_{s}^{2}}+\frac{\alpha_{t|s}^{2}}{\sigma_{t|s}^{2}}\big{)}^{-1}$ and ${\bm{\mu}}_{t\to s}=\sigma_{t\to s}^{2}\big{(}\frac{\alpha_{t|s}}{\sigma_{t|s}^{2}}{\bm{z}}_{t}+\frac{\alpha_{s}}{\sigma_{s}^{2}}{\bm{x}}\big{)}$ as given by (Kingma et al., 2021). In (Song et al., 2021b) it was shown that the optimal solution under a diffusion objective is to learn $\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$, i.e. the expectation over all data given the noisy observation ${\bm{z}}_{t}$. One than iteratively samples for $t=1,1-1/N,\ldots,1/N$ and $s=t-1/N$ starting from ${\bm{z}}_{1}\sim\mathcal{N}(0,1)$. Although the amount of steps required for sampling depends on the data distribution, empirically generative processes for problems such as image generation use hundreds of iterations making diffusion models one of the most resource consuming models to use (Luccioni et al., 2023).
Consistency Models
In contrast, consistency models (Song et al., 2023; Song & Dhariwal, 2023) aim to learn a direct mapping from noise to data. Consistency models are constrained to predict ${\bm{x}}=f({\bm{z}}_{0},0)$, and are further trained by learning to be consistent, minimizing:
$||f({\bm{z}}_{t},t)-\operatorname{nograd}(f({\bm{z}}_{s},s))||,$ (3)
where ${\bm{z}}_{s}=\alpha_{s}{\bm{x}}+\sigma_{s}{\bm{{\epsilon}}}$ and ${\bm{z}}_{t}=\alpha_{t}{\bm{x}}+\sigma_{t}{\bm{{\epsilon}}}$, (note both use the same ${\bm{{\epsilon}}}$) and $s$ is closer to the data meaning $s<t$. When (or if) a consistency model succeeds, the trained model solves for the probability ODE path along time. When successful, the resulting model predicts the same ${\bm{x}}$ along the entire trajectory. At initialization it will be easiest for the model to learn $f$ near zero, because $f$ is defined as an identity function at $t=0$. Throughout training, the model will propagate the end-point of the trajectory further and further to $t=1$. In our own experience, training consistency models is much more difficult than diffusion models.
Consistency Training and Distillation
Consistency Models come in two flavours: Consistency Training (CT) and Consistency Distillation (CD). In the paragraph before, ${\bm{z}}_{s}$ was given by the data which would be the case for CT. Alternatively, one might use a pretrained diffusion model to take a probability flow ODE step (for instance with DDIM). Calling this pretrained model the teacher, the objective for CD can be described by:
$||f({\bm{z}}_{t},t)-\operatorname{nograd}(f(\operatorname{DDIM}_{t\to s}({\bm{x}}_{\mathrm{teacher}},{\bm{z}}_{t}),s))||,$ (4)
where DDIM now defines ${\bm{z}}_{s}$ given the current ${\bm{z}}_{t}$ and (possibly an estimate of) ${\bm{x}}$.
An important hyperparameter in consistency models is the gap between the model evaluations at $t$ and $s$. For CT large gaps incurs a bias, but the solutions are propagated through diffusion time more quickly. On the other hand, when $s\to t$ the bias tends to zero but it takes much longer to propagate information through diffusion time. In practice a step schedule $N(\cdot)$ is used to anneal the step size $t-s=1/N(\cdot)$ over the course of training.
DDIM Sampler
The DDIM sampler is a linearization of the probability flow ODE that is often used in diffusion models. In a variance preserving setting, it is given by:
${\bm{z}}_{s}=\operatorname{DDIM}_{t\to s}({\bm{x}},{\bm{z}}_{t})=\alpha_{s}{\bm{x}}+(\sigma_{s}/\sigma_{t})({\bm{z}}_{t}-\alpha_{t}{\bm{x}})$ (5)
In addition to being a sampling method, the $\operatorname{DDIM}$ equation will also prove to be a useful tool to construct an algorithm for our multistep diffusion models.
Another helpful equations is the inverse of DDIM (Salimans & Ho, 2022), originally proposed to find a natural way parameterize a student diffusion model when a teacher defines the sampling procedure in terms of ${\bm{z}}_{t}$ to ${\bm{z}}_{s}$. The equation takes in ${\bm{z}}_{t}$ and ${\bm{z}}_{s}$, and produces ${\bm{x}}$ for which $\operatorname{DDIM}_{t\to s}({\bm{x}},{\bm{z}}_{t})={\bm{z}}_{s}$. It can be derived by rearranging terms from the $\operatorname{DDIM}$ equation:
${\bm{x}}=\operatorname{invDDIM}_{t\to s}({\bm{z}}_{s},{\bm{z}}_{t})=\frac{{\bm{z}}_{s}-\frac{\sigma_{s}}{\sigma_{t}}{\bm{z}}_{t}}{\alpha_{s}-\alpha_{t}\frac{\sigma_{s}}{\sigma_{t}}}.$ (6)
Figure 2: Qualititative comparison between a multistep consistency and diffusion model. Top: ours, samples from aDDIM distilled 16-step concistency model (3.2 secs). Bottom: generated samples usign a 100-step DDIM diffusion model (39 secs). Both models use the same initial noise.
3 Multistep Consistency Models
In this section we describe multi-step consistency models. First we explain the main algorithm, for both consistency training and distillation. Furthermore, we show that multi-step consistency converges to a standard diffusion training in the limit. Finally, we develop a deterministic sampler named aDDIM that corrects for the missing variance problem in DDIM.
Algorithm 1 Multistep Consistency Models
  Sample ${\bm{x}}\sim p_{\mathrm{data}}$, ${\bm{{\epsilon}}}\sim\mathcal{N}(0,{\mathbf{I}})$, train iteration $i$
  $T_{\mathrm{step}}=\operatorname{round}(N(i)/\mathrm{steps})$
  $\mathrm{step}\sim\mathcal{U}(0,\mathrm{steps}-1)$, $n_{rel}\sim\mathcal{U}(1,T_{\mathrm{step}})$
  $t_{\mathrm{step}}=\mathrm{step}/\mathrm{steps}$
  ${\bm{x}}_{\mathrm{teacher}}=\begin{cases}{\bm{x}}&\text{ if training}\\ f_{\mathrm{teacher}}({\bm{z}}_{t},t)&\text{ if distillation}\\ \end{cases}$
  $x_{\mathrm{var}}=||{\bm{x}}_{\mathrm{teacher}}-{\bm{x}}||^{2}/d$
  $t=t_{\mathrm{step}}+n_{rel}/T$   and   $s=t-1/T$
  ${\bm{z}}_{t}=\alpha_{t}{\bm{x}}+\sigma_{t}{\bm{{\epsilon}}}$
  ${\bm{z}}_{s}=\operatorname{aDDIM}_{t\to s}({\bm{x}}_{\mathrm{teacher}},{\bm{z}}_{t},x_{\mathrm{var}})$
  $\hat{{\bm{x}}}_{\mathrm{ref}}=\operatorname{nograd}(f({\bm{z}}_{s},s))$
  $\hat{{\bm{x}}}=f({\bm{z}}_{t},t)$
  $\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}=\operatorname{DDIM}_{s\to t_{\mathrm{step}}}(\hat{{\bm{x}}}_{\mathrm{ref}},{\bm{z}}_{s})$
  $\hat{{\bm{z}}}_{t_{\mathrm{step}}}=\operatorname{DDIM}_{t\to t_{\mathrm{step}}}(\hat{{\bm{x}}},{\bm{z}}_{t})$
  $\hat{{\bm{x}}}_{\mathrm{diff}}=\operatorname{invDDIM}_{t\to t_{\mathrm{step}}}(\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}},{\bm{z}}_{t})-\hat{{\bm{x}}}$
  $L_{t}=w_{t}\cdot||\hat{{\bm{x}}}_{\mathrm{diff}}||$ for instance $w_{t}=\mathrm{SNR}(t)+1$
Algorithm 2 Sampling from Multistep Consistency
  Sample ${\bm{z}}_{T}\sim\mathcal{N}(0,{\mathbf{I}})$
  for $t$ in $(\frac{\mathrm{steps}}{\mathrm{steps}},\ldots,\frac{1}{\mathrm{steps}})$ where $s=t-\frac{1}{\mathrm{steps}}$ do
     ${\bm{z}}_{s}=\operatorname{DDIM}_{t\to s}(f({\bm{z}}_{t},t),{\bm{z}}_{t})$
  end for
3.1 General description
Multistep consistency split up diffusion time into equal segments to simplify the modelling task. Recall that a consistency model must learn to integrate the full ODE integral. This mapping can become very sharp and difficult to learn when it jumps between modes of the target distribution as can be seen in Figure 1. A consistency loss can be seen as an objective that aims to approximate a path integral by minimizing pairwise discrepancies. Multistep consistency generalizes this approach by breaking up the integral into multiple segments. Originally, consistency runs until time-step 00, evaluated at some time $t>0$. A consistency model should now learn to integrate the DDIM path until 00 and predict the corresponding ${\bm{x}}$. Instead, we can generalize the consistency loss to targets $z_{t_{\mathrm{step}}}$ instead of ${\bm{x}}$ ($\approx{\bm{z}}_{0}$). It turns out that the DDIM equation can be used to operate on ${\bm{z}}_{t_{\mathrm{step}}}$ for different times $t_{\mathrm{step}}$, which allows us to express the multi-step consistency loss as:
$||\operatorname{DDIM}_{t\to t_{\mathrm{step}}}(f({\bm{z}}_{t},t),{\bm{z}}_{t})-\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}||,$ (7)
where $\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}=\operatorname{DDIM}_{s\to t_{\mathrm{step}}}(\operatorname{nograd}f({\bm{z}}_{s},s))$ and where the teaching step ${\bm{z}}_{s}=\operatorname{aDDIM}_{t\to s}(x,{\bm{z}}_{t})$.
For now it suffices to think $\operatorname{aDDIM}$ as $\operatorname{DDIM}$. It will be described in detail in section 3.2. In fact, one can drop-in any deterministic sampler in place of $\operatorname{aDDIM}$ in the case of distillation.
A model can be trained on directly on this loss in $z$ space, however make the loss more interpretable and relate it more closely to standard diffusion, we re-parametrize the loss to $x$-space using:
$||\hat{{\bm{x}}}_{\mathrm{diff}}||=||f({\bm{z}}_{t},t)-\operatorname{invDDIM}_{t\to t_{\mathrm{step}}}(\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}},{\bm{z}}_{t})||.$ (8)
This allows the usage of existing losses from diffusion literature, where we have opted for $v$-loss (equivalent to $\mathrm{SNR}+1$ weighting) because of its prior success in distillation (Salimans & Ho, 2022).
As noted in (Song et al., 2023), consistency in itself is not sufficient to distill a path (always predicting 00 is consistent) and one needs to ensure that the model cannot collapse to these degenerate solutions. Indeed, in our specification observe that that $\operatorname{DDIM}_{t_{\mathrm{step}}\to t_{\mathrm{step}}}({\bm{z}}_{t_{\mathrm{step}}},\hat{\bm{x}})={\bm{z}}_{t_{{\mathrm{step}}}}$ and that for the final step in each segment $s=t_{\mathrm{step}}$. As such, the loss of the final step cannot be degenerate and is equal to:
$||f({\bm{z}}_{t},t)-\operatorname{invDDIM}_{t_{\mathrm{step}}\to t}({\bm{z}}_{t_{\mathrm{step}}})||.$ (9)
Many-step CT is equivalent to Diffusion training
Consistency training learns to integrate the probability flow through time, whereas standard diffusion models learn a path guided by an expectation $\hat{{\bm{x}}}=\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$ that necessarily has to change over time for non-trivial distributions. There are two simple reasons that for many steps, Multistep CT converges to a diffusion model. 1) At the beginning of a step (specifically $t={t_{\mathrm{step}}}+\frac{1}{T}$) the objectives are identical. Secondly, 2) when the number of steps equals the total number of steps $T$, then every step is equal to the diffusion objective. This can be observed by studying Algorithm 1: let $t=t_{\mathrm{step}}+1/T$. For consistency training, aDDIM reduces to DDIM and observe that in this case $s=t_{\mathrm{step}}$. Hence, under a well-defined model $f$ (such as a $v$-prediction one) $\operatorname{DDIM}_{s\to t_{\mathrm{step}}}$ does nothing and simply produces $\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}={\bm{z}}_{s}$. Also observe that $\hat{{\bm{z}}}_{t_{\mathrm{step}}}=\hat{{\bm{z}}}_{s}$. Further simplification yields:
$\begin{split}w(t)||{\bm{x}}_{\mathrm{diff}}||&=\operatorname{invDDIM}_{t\to s}(\hat{{\bm{z}}}_{s},{\bm{z}}_{t})-\hat{{\bm{x}}}\\ &=w(t)||{\bm{x}}-\hat{{\bm{x}}}||\\ \end{split}$ (10)
Where $||{\bm{x}}-\hat{{\bm{x}}}||$ is the distance between the true datapoint and the model prediction weighted by $w(t)$, which is typical for standard diffusion. Interestingly, in (Song & Dhariwal, 2023) it was found that Euclidean ($\ell_{2}$) distances typically work better than for consistency models than the more usual squared Euclidean distances ($\ell_{2}$ squared). We follow their approach because it tended to work better especially for smaller steps, which is a deviation from standard diffusion. Because multistep consistency models tend towards diffusion models, we can state two important hypotheses:
  1. 1.
    Finetuning Multistep CMs from a pretrained diffusion checkpoint will lead to quicker and more stable convergence.
  2. 2.
    As the number of steps increases, Multistep CMs will rival diffusion model performance, giving a direct trade-off between sample quality and duration.
Figure 3: Another qualititative comparison between a multistep consistency and diffusion model. Top: ours, samples from aDDIM distilled 16-step concistency model (3.2 secs). Bottom: generated samples using a 100-step DDIM diffusion model (39 secs). Both models use the same initial noise.
What about training in continuous time?
Diffusion models can be easily trained in continuous time by sampling $t\sim\mathcal{U}(0,1)$, but in Algorithm 1 we have taken the trouble to define $t$ as a discrete grid on $[0,1]$. One might ask, why not let $t$ be continuously valued. This is certainly possible, if the model $f$ would take in an additional conditioning signal to denote in which step it is. This is important because its prediction has to discontinuously change between $t\geq t_{\mathrm{step}}$ (this step) and $t<t_{\mathrm{step}}$ (the next step). In practice, we often train Multistep Consistency Models starting from pre-trained with standard diffusion models, and so having the same interface to the model is simpler. In early experiments we did find this approach to work comparably.
Algorithm 3 Generating Samples with aDDIM
  Sample ${\bm{z}}_{T}\sim\mathcal{N}(0,{\mathbf{I}})$, choose hyperparam $10<\eta<$ 1
  For all $t$, precompute $x_{\mathrm{var},t}=\eta||{\bm{x}}-\hat{{\bm{x}}}({\bm{z}}_{t})||^{2}/d$, or set $x_{\mathrm{var},t}=0.1/(2+\alpha^{2}_{t}/\sigma^{2}_{t})$.
  for $t$ in $(\frac{T}{T},\ldots,\frac{1}{T})$ where $s=t-1/T$ do
     $\hat{{\bm{x}}}=f({\bm{z}}_{t},t)$
     $\hat{{\bm{{\epsilon}}}}=({\bm{z}}_{t}-\alpha_{t}\hat{{\bm{x}}})/\sigma_{t}$
     $z_{s,\mathrm{var}}=(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})^{2}\cdot x_{\mathrm{var},t}$
     ${\bm{z}}_{s}=\alpha_{s}\hat{{\bm{x}}}+\sqrt{\sigma_{s}^{2}+(d/||\hat{{\bm{{\epsilon}}}}||^{2})z_{s,\mathrm{var}}}\cdot\hat{{\bm{{\epsilon}}}}$
  end for
3.2 The Adjusted DDIM (aDDIM) sampler.
Popular methods for distilling diffusion models, including the method we propose here, rely on deterministic sampling through numerical integration of the probability flow ODE. In practice, numerical integration of this ODE in a finite number of steps incurs error. For the DDIM integrator (Song et al., 2021a) used for distilling diffusion models in both consistency distillation (Song et al., 2023) and progressive distillation (Salimans & Ho, 2022; Meng et al., 2022) this integration error causes samples to become blurry. To see this quantitatively, consider a hypothetical perfect sampler that first samples ${\bm{x}}^{*}\sim p({\bm{x}}|{\bm{z}}_{t})$, and then samples ${\bm{z}}_{s}$ using
${\bm{z}}^{*}_{s}=\alpha_{s}{\bm{x}}^{*}+\sigma_{s}\frac{{\bm{z}}_{t}-\alpha_{t}{\bm{x}}^{*}}{\sigma_{t}}=(\alpha_{s}-\frac{\alpha_{t}\sigma_{s}}{\sigma_{t}}){\bm{x}}^{*}+\frac{\sigma_{s}}{\sigma_{t}}{\bm{z}}_{t}.$
If the initial ${\bm{z}}_{t}$ is from the correct distribution $p({\bm{z}}_{t})$, the sampled ${\bm{z}}^{*}_{s}$ would then also be exactly correct. Instead, the DDIM integrator uses
${\bm{z}}^{\text{DDIM}}_{s}=(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})\hat{{\bm{x}}}+(\sigma_{s}/\sigma_{t}){\bm{z}}_{t},$
with model prediction $\hat{{\bm{x}}}$. If $\hat{{\bm{x}}}=\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$, we then have that
$||{\bm{z}}^{*}_{s}||^{2}-||{\bm{z}}^{\text{DDIM}}_{s}||^{2}=\text{trace}(\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}]),$
where $\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}]$ is the conditional variance of ${\bm{z}}_{s}$ given by
$\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}]=(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})^{2}\cdot\mathrm{Var}[{\bm{x}}|{\bm{z}}_{t}],$
and where $\mathrm{Var}[{\bm{x}}|{\bm{z}}_{t}]$ in turn is the variance of $p({\bm{x}}|{\bm{z}}_{t})$.
The norm of the DDIM iterates is thus too small, reflecting the lack of noise addition in the sampling algorithm. Alternatively, we could say that the model prediction $\hat{{\bm{x}}}\approx\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$ is too smooth. Second-order integrators like the Heun sampler proposed by Karras et al. (2022) aim to remove this bias by using the prediction at $s$ in addition to that at $t$, but in practice these methods did not significantly improve sample quality in our experiments.
Currently, the best sample quality is achieved with stochastic samplers, which can be tuned to add exactly enough noise to undo the oversmoothing caused by numerical integration. However, current distillation methods are not well suited to distilling these stochastic samplers directly. Here we therefore propose a new deterministic sampler that aims to achieve the norm increasing effect of noise addition in a deterministic way. It turns out we can do this by making a simple adjustment to the DDIM sampler, and we therefore call our new method Adjusted DDIM (aDDIM). Our modification is heuristic and is not more theoretically justified than the original DDIM sampler. However, empirically we find aDDIM to work very well leading to improved FID scores.
Instead of adding noise to our sampled ${\bm{z}}_{s}$, we simply increase the contribution of our deterministic estimate of the noise $\hat{{\bm{{\epsilon}}}}=({\bm{z}}_{t}-\alpha_{t}\hat{{\bm{x}}})/\sigma_{t}$. Assuming that $\hat{{\bm{x}}}$ and $\hat{{\bm{{\epsilon}}}}$ are orthogonal, we achieve the correct norm for our sampling iterates using:
${\bm{z}}^{\text{aDDIM}}_{s}=\alpha_{s}\hat{{\bm{x}}}+\sqrt{\sigma_{s}^{2}+\text{tr}(\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}])/||\hat{{\bm{{\epsilon}}}}||^{2}}\cdot\hat{{\bm{{\epsilon}}}}.$ (11)
In practice, we can estimate $\text{tr}(\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}])=(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})^{2}\cdot\text{tr}(\mathrm{Var}[{\bm{x}}|{\bm{z}}_{t}])$ empirically on the data by computing beforehand $\text{tr}(\mathrm{Var}[{\bm{x}}|{\bm{z}}_{t}])=\eta||\hat{{\bm{x}}}({\bm{z}}_{t})-{\bm{x}}||^{2}$ for all relevant timesteps $t$. Here $\eta$ is a hyperparameter which we set to $0.75$. Alternatively, we obtain equally good results by approximating the posterior variance analytically with $\text{tr}(\mathrm{Var}[{\bm{x}}|{\bm{z}}_{t}])/d=0.1/(2+\alpha^{2}_{t}/\sigma^{2}_{t})$, for data dimension $d$, which can be interpreted as $10\%$ of the posterior variance of ${\bm{x}}$ if its prior was factorized Gaussian with variance of $0.5$. In either case, note that $\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}]$ vanishes as $s\rightarrow t$: in the many-step limit the aDDIM update thus becomes identical to the original DDIM update. For a complete description see Algorithm 3.
4 Related Work
Existing works closest to ours are Consistency Models (Song et al., 2023; Song & Dhariwal, 2023) and TRACT (Berthelot et al., 2023). Compared to consistency models, we propose to operate on multiple stages, which simplifies the modelling task and improves performance significantly. On the other hand, TRACT limits itself to distillation and uses the self-evaluation from consistency models to distill models over multiple stages. The stages are progressively reduced to either one or two stages and thus steps. The end-goal of TRACT is again to sample in either one or two steps, whereas we believe better results can be obtained by optimizing for a slightly larger number of steps. We show that this more conservative target, in combination with our improved sampler and annealed schedule, leads to significant improvements in terms of image quality that closes the gap between sample quality of standard diffusion and low-step diffusion-inspired approaches.
Earlier, DDIM (Song et al., 2021a) showed that deterministic samplers degrade more gracefully than the stochastic sampler used by Ho et al. (2020) when limiting the number of sampling steps. Karras et al. (2022) proposed a second order Heun sampler to reduce the number of steps (and function evaluations), while Jolicoeur-Martineau et al. (2021) studied different SDE integrators to reduce function evaluations. Zheng et al. (2023) use specialized architectures to distill the ODE trajectory from a pre-created noise-sample pair dataset. Progressive Distillation (Salimans & Ho, 2022; Meng et al., 2022) distills diffusion models in stages, which limits the number of model evaluations during training while exponentially reducing the required number of sampling steps with the number stages. Luo et al. (2023) distill the knowledge from the diffusion model into a single-step model.
Other methods inspired by diffusion such as Rectified Flows (Liu et al., 2023) and Flow Matching (Lipman et al., 2023) have also tried to reduce sampling times. In practice however, flow matching and rectified flows are generally used to map to a standard normal distribution and reduce to standard diffusion. As a consequence, on its own they still require many evaluation steps. In Rectified Flows, a distillation approach is proposed that does reduce sampling steps more significantly, but this comes at the expense of sample quality.
Table 1: Imagenet performance with multistep consistency training (CT) and consistency distillation (CD), started from a pretrained diffusion model. A baseline with the aDDIM sampler on the base model is included.
ImageNet64 ImageNet128
Steps Train Distill Train Distill
Base Consistency Model 1 7.2 4.3 16.0 8.5
MultiStep CM (ours) 2 2.7 2.0 6.0 3.1
MultiStep CM (ours) 4 1.8 1.7 4.0 2.4
MultiStep CM (ours) 8 1.5 1.6 3.3 2.1
MultiStep CM (ours) 16 1.5 1.5 3.4 2.0
Diffusion (aDDIM) 512 1.5 2.2
Large Consistency Model 1 6.4 3.2 14.5 7.0
MultiStep CM (ours) 2 2.3 1.9 4.2 3.1
MultiStep CM (ours) 4 1.6 1.6 2.7 2.3
MultiStep CM (ours) 8 1.5 1.4 2.2 2.1
MultiStep CM (ours) 16 1.6 1.4 2.3 2.0
Diffusion (aDDIM) 512 1.4 2.2
Table 2: Ablation of CD on Image128 with and without annealing the steps on ImageNet128.
Steps (64 $\to$ 1280) (step = 128) (step = 256) (step = 1024)
1 7.0 8.8 7.6 10.8
2 3.1 5.3 3.6 3.8
4 2.3 5.0 3.5 2.6
8 2.1 4.9 3.2 2.2
Table 3: Comparison between Progressive Distillation (Salimans & Ho, 2022) and CT and CD on ImageNet64 on the base model variant. Reimplementation of PD with modern optimizations.
Steps CT (ours) CD (ours) PD
1 7.2 4.3 10.7
2 2.7 2.0 4.7
4 1.8 1.7 2.4
8 1.5 1.6 1.8
5 Experiments
Our experiments focus on a quantitative comparison using the FID score on ImageNet as well as a qualitative assessment on large scale Text-to-Image models. These experiments should make our approach comparable to existing academic work while also giving insight in how multi-step distillation works at scale.
5.1 Quantitative Evaluation on ImageNet
For our ImageNet experiments we trained diffusion models on ImageNet64 and ImageNet128 in a base and large variant. We initialize the consistency models from the pre-trained diffusion model weights which we found to greatly increase robustness and convergence. Both consistency training and distillation are used. Classifier Free Guidance (Ho & Salimans, 2022) was used only on the base ImageNet128 experiments. For all other experiments we did not use guidance because it did not significantly improve the FID scores of the diffusion model. All consistency models are trained for $200,000$ steps with a batch size of $2048$ and an step schedule that anneals from $64$ to $1280$ in $100.000$ steps with an exponential schedule.
In Table 1 it can be seen that as the multistep count increases from a single consistency step, the performance considerably improves. For instance, on the ImageNet64 Base model it improves from 7.2 for one-step, to 2.7 and further to 1.8 for two and four-steps respectively. In there are generally two patterns we observe: As the steps increase, performance improves. This validates our hypothesis that more steps give a helpful trade-off between sample quality and speed. It is very pleasant that this happens very early: even on a complicated dataset such as ImageNet128, our base model variant is able achieve 2.1 FID in 8 steps, when consistency distilling.
To draw a direct comparison between Progressive Distillation (PD) (Salimans & Ho, 2022) and our approaches, we reimplement PD using aDDIM and we use same base architecture, as reported in Table 3. With our improvements, PD can attain better performance than previously reported in literature. However, compared to MultiStep CT and CD it starts to degrade in sample quality at low step counts. For instance, a 4-step PD model attains an FID of 2.4 whereas CD achieves 1.7.
Further we are ablating whether annealing the step schedule is important to attain good performance. As can be seen in Table 2, it is especially important for low multistep models to anneal the schedule. In these experiments, annealing always achieves better performance than tests with constant steps at $128,256,1024$. As more multisteps are taken, the importance of the annealing schedule is less important.
Figure 4: Yet another qualititative comparison between a multistep consistency and diffusion model. Top: ours, this time using a standard DDIM distilled 16-step concistency model (3.2 secs). Bottom: generated samples using a 100-step DDIM diffusion model (39 secs). Both models use the same initial noise.
Table 4: Literature Comparison on ImageNet64 and 128.
Method NFE FID
Imagenet 64 x 64
DDIM (Song et al., 2021a) 10 18.7
DFNO (LPIPS) (Zheng et al., 2023) 1 7.83
TRACT (Berthelot et al., 2023) 1 7.43
2 4.97
4 2.93
8 2.41
Diff-Instruct 1 5.57
PD (Salimans & Ho, 2022) 1 10.7
     (reimpl. with aDDIM) 2 4.7
4 2.4
8 1.7
PD Stochastic (Meng et al., 2022) 1 18.5
2 5.81
4 2.24
8 2.31
CD (LPIPS) (Song et al., 2023) 1 6.20
2 4.70
3 4.32
PD (LPIPS) (Song et al., 2023) 1 7.88
2 5.74
3 4.92
iCT-deep (Song & Dhariwal, 2023) 1 3.25
iCT-deep 2 2.77
MultiStep-CT (ours) 2 2.3
4 1.6
8 1.5
MultiStep-CD (ours) 1 3.2
2 1.9
4 1.6
8 1.4
Imagenet 128 x 128
VDM++ (Kingma & Gao, 2023) 512 1.75
PD (Salimans & Ho, 2022) 2 8.0
     (reimpl. with aDDIM) 4 3.8
8 2.5
MultiStep-CT (ours) 2 4.2
4 2.7
8 2.2
MultiStep-CD (ours) 2 3.1
4 2.3
8 2.1
Literature Comparison
Compared to existing works in literature, we achieve SOTA FID scores in both ImageNet64 on 2-step, 4-step and 8-step generation. Interestingly, we achieve approximately the same performance using single step CD compared to iCT-deep (Song & Dhariwal, 2023), which achieves this result using direct consistency training. Since direct training has been empirically shown to be a more difficult task, one could conclude that some of our hyperparameter choices may still be suboptimal in the extreme low-step regime. Conversely, this may also mean that multistep consistency is less sensitive to hyperparameter choices.
In addition, we compare on ImageNet128 to our reimplementation of Progressive Distillation. Unfortunately, ImageNet128 has not been widely adopted as a few-step benchmark, possibly because a working deterministic sampler has been missing until this point. For reference we also provide the recent result from (Kingma & Gao, 2023). Further, with these results we hope to put ImageNet128 on the map for few-step diffusion model evaluation.
5.2 Qualitative Evaluation on Text to Image modelling
In addition to the quantitative analysis on ImageNet, we study the effects on a text-to-image model by directly comparing samples. We first train a 20B parameter diffusion model on text-to-image pairs with a T5 XXL paper following (Saharia et al., 2022) for 1.3 million steps. Then, we distill a 16-step consistency model using the DDIM sampler. In Figure 2 and 3 we compare samples from our 16-step CD aDDIM distilled model to the original 100-step DDIM sampler. Because the random seed is shared we can easily compare the samples between these models, and we can see that there are generally minor differences. In our own experience, we often find certain details more precise, at a slight cost of overall construction. Another comparison in Figure 4 shows the difference between a DDIM distilled model (equivalent to $\eta=0$ in aDDIM) and the standard DDIM sampler. Again we see many similarities when sharing the same initial random seed.
6 Conclusions
In conclusion, this paper presents Multistep Consistency Models, a simple unification between Consistency Models (Song et al., 2023) and TRACT (Berthelot et al., 2023) that closes the performance gap between standard diffusion and few-step sampling. Multistep Consistency gives a direct trade-off between sample quality and speed, achieving performance comparable to standard diffusion in as little as eight steps.